# FedEFC: Federated Learning Using Enhanced Forward Correction Against Noisy Labels

### v1

@@ 1 Introduction
Federated Learning (FL) is a powerful paradigm for distributed learning that enables the training of a high-performing global model without requiring the aggregation or centralization of locally stored data [13]. While FL provides strong privacy guarantees by keeping client data decentralized, this non-centralized nature makes the model highly sensitive to the underlying data distribution among clients. In particular, the challenge of convergence under heterogeneous data distributions has been extensively studied [11, 10, 6, 1].
Recently, a growing body of research [23, 22, 3, 24] has investigated the impact of noisy datasets in addition to data heterogeneity. The adverse effects of label noise—whether caused by natural annotation errors or adversarial attacks—are often more pronounced in FL than in centralized learning (CL) due to the decentralized nature of training and the aggregation of corrupted model updates. Moreover, the lack of direct access to client datasets significantly limits the applicability of conventional noise mitigation techniques commonly used in CL, necessitating novel approaches tailored to the FL setting.
@@
  • We provide a theoretical proof demonstrating that the enhanced forward correction enables each client to achieve the comparable training effectiveness as if learning from entirely clean data, despite the presence of noisy labels. (Sec. 3).
  • Experimental results confirm that FedEFC outperforms existing FL techniques, demonstrating its robustness in mitigating the impact of noisy labels and ensuring reliable model performance (Sec. 4). To achieve this, we introduce sparsity, a measure that quantifies the degree of asymmetric label noise, into our noisy label synthesis process [15]. Additionally, we leverage the Dirichlet distribution and Bernoulli distribution to systematically allocate data in a heterogeneous manner, ensuring a realistic simulation of non-IID conditions in FL. (Sec. 2).
@@ 1.1 Related Works
In [24], count matrix has also been integrated into FL for label correction. However, this approach has notable limitations, as it does not explicitly account for heterogeneous data distributions across clients and depends on a pretrained model for reliable performance, which may not always be feasible in real-world FL scenarios. Though, since adapting confident learning in FL offers a meaningful baseline for comparison with FedEFC, we modify it to ensure a fair and consistent evaluation within FL setting.
@@
FL algorithms for non-IID settings : In real-world FL scenarios, data is typically distributed in a non-IID manner, introducing significant challenges for model training [5, 25]. Various strategies have been proposed to enhance FL performance under heterogeneous settings. For instance, FedProx [10] introduces minor modifications to FedAvg to achieve more robust convergence when training on non-IID data. FedDyn [1] dynamically aligns the local optimization objectives with the global loss function, ensuring more stable updates across clients. Ditto [11] enhances data robustness and fairness, improving individual client performance while maintaining overall model consistency. In this work, we consider these methods as baseline algorithms for comparison with FedEFC, evaluating its effectiveness in mitigating noisy labels and handling heterogeneous FL settings.
FL algorithms against noisy dataset : Several research efforts have explored ways to resolve the noise labels present in local datasets of FL. In [22], FedCorr is designed to cope with noisy data challenges. On the other hand, the framework requires the assumption that certain clients have entirely clean data to ensure improvement. RHFL [3] proposes a robust noise loss function for noisy labels under non-IID whereas public data is necessary to utilize the proposed algorithm. Although RoFL [23] effectively leverages similarity-based learning to mitigate the issue of asymmetric noisy labels, it requires additional information on feature data that may pose a potential risk to privacy and communication bottleneck. Of the approaches considered, FedCorr is compared with the proposed method as it serves as a more suitable baseline for a fair evaluation. Unlike methods relying on additional public data (RHFL) or feature-based similarity learning (RoFL), FedCorr better aligns with the constraints and challenges of real-world FL scenarios.
Figure 1: Overview of FedEFC framework. The scheme consists of two phases: (1) determining the prestopping point and (2) refining loss correction. In Phase 1, the centralized server tracks client training accuracies to identify the prestopping point where model parameters are near-optimal. In Phase 2, each client updates its loss function using enhanced forward correction, guiding global model parameters toward their optimal configuration in the clean data space.
@@ 2.1 Data Allocation
Specifically, the Bernoulli probability $p$ determines whether examples of a specific label $i$ are allocated to client $k$. The Bernoulli distribution is represented by the indicator $\mathbb{I}(k,i)\sim p$, which takes the value $1$ if label $i$ is allocated to client $k$, and 00 otherwise. Once the Bernoulli distribution is established, the number of examples assigned to clients with $\mathbb{I}(\cdot,i)=1$ follows a Dirichlet distribution parameterized by $\alpha_{\text{dir}}>0$. Therefore, the degree of non-IID data allocation is controlled by the parameters $p$ and $\alpha_{\text{dir}}$.
@@ 2.2 Noise Generation
The asymmetric noise flip more closely resembles real-world scenarios than the symmetric case when generating synthetic noise. To generate asymmetric noise, the sparsity mentioned in [15], [20] is utilized. The amount of noise $\rho$ and sparsity $\zeta$ need to be considered for noise generation. $\rho$ is the fraction of samples whose labels are flipped while $\zeta$ represents the proportion of uncorrupted labels, excluding the true label. For example, if all labels are flipped, $\zeta=0$. In contrast, $\zeta=1$ indicates that the dataset contains only clean labels. Thus, high sparsity represents greater imbalance in noisy labels.
@@ 3 Proposed Method
In this section, we propose FedEFC for mitigating the impact of noisy labels in non-IID FL. FedEFC consists of two phases: (1) determining a prestopping point and (2) applying loss correction. In Phase 1, the centralized server estimates the accuracy of the global model by aggregating the training accuracies of local clients. During Phase 1, each client transmits its measured accuracy to the centralized server, which monitors accuracy variations to determine the prestopping point. Once this point is reached, the server notifies the clients that Phase 1 is complete and that no further accuracy updates will be transmitted in subsequent rounds. In Phase 2, participating clients generate a noise transition matrix by analyzing the allocated dataset and apply loss correction. Detailed descriptions of each phase are provided below.
@@ 3.1 Finding the Prestopping Point
where $A_{k}(w_{g}^{t})$ is training accuracy of client $k$ with the global model based on its own dataset. The training set contains both clean and noisy labels, which differs from the original heuristic validation approach that relies on a clean set [19]. A threshold $\gamma_{\text{thr}}$ is defined to quantify the instability in learning. In addition, a patience parameter $\tau_{p}$ is employed to track previous accuracy history. The accuracy measured in the current round, $A(t)$, is compared to that of the previous round, $A(t-1)$. If $A(t)<A(t-1)$, the patience parameter $\tau_{p}$ is incremented by 1; otherwise, $\tau_{p}$ is reset to 00. When the accuracy does not improve for $\gamma_{\text{thr}}$ consecutive rounds (i.e., $\tau_{p}=\gamma_{\text{thr}}$), the current round is regarded as the prestopping point, denoted as $T_{e}$. At the end of Phase 1, the centralized server instructs the clients to apply loss correction for every round after $T_{e}$. Empirical verification of the estimated accuracy, aggregated from each client’s training performance, is provided in the left panel of Fig. 2. In this panel, the estimated accuracy exhibits a sudden decline, attributed to the presence of noisy labels.
Figure 2: Left: Test accuracy and estimated accuracy used to determine the prestopping point. Right: Cosine similarity between the real noise matrix and the noise matrices estimated by the pretrained model and the global model in training. Experiments are conducted on the CIFAR-10 with the following settings: $\alpha_{\text{dir}}=10.0$, $p=0.5$, $\rho=0.2$, $\zeta=0.8$, and $T_{e}=54$.
@@ 3.2.1 Generating the Noise Transition Matrix
Each client acquires the noise transition matrix as follows. For client $k$, the global model after round $T_{e}$ predicts the true label as
$j=\arg\max_{\tilde{y}\in\mathcal{C}}\hat{p}(\tilde{y};\bm{x}^{k},w^{t}_{g}),$ (3)
where $\hat{p}_{k}(\cdot)$ denotes the estimated probability for client $k$. We define the set of examples with true label $y=j$ and observed label $\tilde{y}=i$ as
@@
=$\quad\quad=\Bigl{\{}\bm{x}^{k}\in\mathbf{X}_{\tilde{y}^{k}=i}:\hat{p}_{k}(\tilde{y}^{k}=j;\bm{x}^{k},w^{t}_{g})\geq\tau_{j}\Bigr{\}},$ (4)
where the set $\mathbf{X}_{\tilde{y}^{k}=i}$ represents all data samples with observed label $i$ in the dataset $\mathcal{D}_{k}$ of client $k$. The example set $\mathbf{X}_{\tilde{y}^{k}=i}$ satisfies
$\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\subseteq\mathbf{X}_{\tilde{y}^{k}=i}\subseteq\mathcal{D}_{k}.$ (5)
The threshold $\tau_{j}$ serves as the criterion for counting examples and is defined as
$\tau_{j}=\frac{1}{\left|\mathbf{X}_{\tilde{y}^{k}=j}\right|}\sum_{\bm{x}^{k}\in\mathbf{X}_{\tilde{y}^{k}=j}}\hat{p}_{k}(\tilde{y}=j;\bm{x}^{k},w^{t}_{g}),$ (6)
Threshold $\tau_{j}$ is the mean estimation probabilities for data labeled as $j$. Using the example set $\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}$, the count matrix for client $k$ is defined as
$\bm{C}_{\tilde{y}^{k}=i,y^{k}=j}=\left|\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\right|.$ (7)
For simplicity, the notation is rewritten as
$\bm{C}_{\tilde{y}^{k}=i,y^{k}=j}\coloneqq\bm{C}^{k}_{\tilde{y}=i,y=j},\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\coloneqq\mathbf{X}^{k}_{\tilde{y}=i,y=j}.$ (8)
The count matrix is constructed to enumerate the number of examples associated with each label pair, with columns representing true labels and rows indicating observed labels. Fig. 3 illustrates an example of count matrix construction. The noise transition matrix of client $k$ is computed as
@@
$\bm{Q}^{k}_{\tilde{y}=i|y=j}$ represents the $(i,j)$th element of the noise transition matrix $\bm{Q}^{k}_{{\tilde{y}}|{y}}$. This element corresponds to the conditional probability $p(\tilde{y}=i\mid y=j)$, which indicates the probability that the observed noisy label $\tilde{y}$ is $i$ given that the true label $y$ is $j$. Consequently, each client generates its own noise transition matrix $\bm{Q}^{k}_{{\tilde{y}}|{y}}$.
@@ 3.2.2 Applying Loss Correction
Participating clients apply loss correction based on the derived matrix $\bm{Q}_{{\tilde{y}}|{y}}$, following a similar approach to [16], where loss functions are modified. Regarding client $k$ in round $t$, the loss function $\mathcal{L}_{k}(\cdot)$ is given by
$\mathcal{L}_{k}(\bm{x},y;{w}_{k}^{t}).$ (10)
When a client trains with a DNN and applies the softmax function $\sigma_{k}(\cdot)$ in the final layer, the probability of the observed label $\tilde{y}=i$ given input $\bm{x}^{k}$ is
@@
$\bm{h}_{i}^{k}$ denotes the output of the DNN at the final layer corresponding to label $i$. The loss function $\mathcal{L}_{k}$ can be rewritten using cross-entropy as
$\mathcal{L}_{k}(\hat{p}_{k}(\tilde{y}^{k}=i|\bm{x}^{k};w_{k}^{t}))=-\log\hat{p}_{k}(\tilde{y}=i|\bm{x};w_{k}^{t}).$ (13)
To streamline notation, we replace $\bm{x}^{k}$ and $y^{k}$ with $\bm{x}$ and $y$, respectively. Using the noise transition matrix $Q_{\tilde{y}=i|{y}=j}^{k}$, the loss function $\mathcal{L}_{k}^{\textbf{F}}(\cdot)$ is updated as follows:
$\mathcal{L}_{k}^{\textbf{F}}(\hat{p}_{k}(\tilde{y}=i|\bm{x};w_{k}^{t}))=-\log{\hat{p}_{k}(\tilde{y}=i|\bm{x};w_{k}^{t})}$
$=-\log\left\{\sum_{j\in\mathcal{C}}p(\tilde{y}={i}|{y}=j)\cdot\hat{p}_{k}({y}=j|\bm{x};w_{k}^{t})\right\}$ (14)
$=-\log\left\{\sum_{j\in\mathcal{C}}Q^{k}_{\tilde{y}=i|y=j}\cdot\hat{p}_{k}({y}=j|\bm{x};w_{k}^{t})\right\}$ (15)
$=-\log\left\{Q_{\tilde{y}=i|{y}}^{k}\cdot\hat{p}_{k}({y}|\bm{x};w_{k}^{t})\right\}.$ (16)
$Q_{\tilde{y}=i|{y}}^{k}$ denotes the row vector of the noise transition matrix $Q^{k}_{{\tilde{y}}|{y}}$. The enhanced forward correction is obtained by weighting the loss function with the noise transition matrix $Q_{\tilde{y}=i|{y}}^{k}$. Clients then begin transmitting their learning parameters, updated via loss correction, to the centralized server. The entire process is summarized in Algorithm 1.
@@
4:for $t=0$ to $T_{e}$ do
5:     $\mathcal{S}_{t}=Sampling({1,2,...,N})$; $\tau_{p}=0$
6:     for each client $k\in\mathcal{S}_{t}$ in parallel do
7:         $\left(w_{k}^{t+1},A_{k}(w_{g}^{t})\right)$leftarrow Client update($k$, $w_{g}^{t}$)
8:         Compute $A(t)$ through Eq. (2)
9:         if $A(t)>A({t-1})$ then
10:              $\tau_{p}\leftarrow 0$
11:         else
12:              $\tau_{p}\leftarrow\tau_{p}+1$
13:              if $\tau_{p}==\gamma_{\text{thr}}$ then
14:                  return $T_{e}=t$
15:              end if
16:         end if
17:     end for
18:     $w_{g}^{t+1}\leftarrow\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}}{\sum_{u\in\mathcal{S}_{t}}n_{u}}w_{k}^{t+1}$
@@
20:Phase 2: Apply loss correction
21:for $t=T_{e}$ to $T$ do
22:     for each client $k\in\mathcal{S}_{t}$ in parallel do
23:         Compute ${Q}^{k}_{y=j|\tilde{y}=i}$ via Eq. (9)
24:         Update $\mathcal{L}_{k}^{\textbf{F}}(\bm{x},y;w_{k}^{t})$ using Eq. (3.2.2)
25:         $w_{k}^{t+1}$leftarrow Client update($k$, $w_{g}^{t}$, $\mathcal{L}_{k}^{\textbf{F}}(\cdot)$)
26:     end for
27:     $w_{g}^{t}\leftarrow\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}}{\sum_{u\in\mathcal{S}_{t}}n_{u}}w_{k}^{t}$
28:end for

### v3

@@ 1 Introduction
Federated Learning (FL) is a powerful paradigm for distributed learning that enables the training of a high-performing global model without requiring the aggregation or centralization of locally stored data [14]. While FL provides strong privacy guarantees by keeping client data decentralized, this non-centralized nature makes the model highly sensitive to the underlying data distribution among clients. The challenge of convergence under heterogeneous data distributions has been extensively studied [12, 10, 11, 6, 1].
Recently, a growing body of research [25, 24, 3, 23, 26] has investigated the impact of noisy datasets in addition to data heterogeneity. The adverse effects of label noise—whether caused by natural annotation errors—are often more pronounced in FL than in centralized learning (CL) due to the decentralized nature of training and the aggregation of corrupted model updates. Moreover, the lack of direct access to client datasets significantly limits the applicability of conventional noise mitigation methods commonly used in CL, necessitating novel approaches tailored to the FL setting.
@@
  • We provide a theoretical proof demonstrating that the enhanced forward correction enables each client to achieve the comparable training effectiveness as if learning from entirely clean data, despite noisy labels. (Sec. 3).
  • Experimental results confirm that FedEFC outperforms existing FL techniques, demonstrating its robustness in mitigating the impact of noisy labels and ensuring reliable model performance (Sec. 4). To achieve this, we incorporate sparsity [16], a measure that quantifies the degree of asymmetric label noise, into our noisy label synthesis process. Additionally, we leverage the Dirichlet distribution and Bernoulli distribution to systematically allocate data in a heterogeneous manner, ensuring a realistic simulation of non-IID conditions in FL. (Sec. 2).
@@ 1.1 Related Works
In [26], count matrix has also been integrated into FL for label correction. However, this approach has notable limitations, as it does not explicitly account for heterogeneous data distributions across clients and depends on a pretrained model for reliable performance, which may not always be feasible in real-world FL scenarios. Though, since adapting confident learning in FL offers a meaningful baseline for comparison with FedEFC, we modify it to ensure a fair and consistent evaluation within the FL setting.
@@
FL algorithms for non-IID settings: In real-world FL scenarios, data is typically distributed in a non-IID manner, introducing significant challenges for model training [5, 27]. Various strategies have been proposed to enhance FL performance under heterogeneous settings. For instance, FedProx [11] introduces minor modifications to FedAvg to achieve more robust convergence when training on non-IID data. Ditto [10] enhances data robustness and fairness, improving individual client performance while maintaining overall model consistency. In this work, we consider these methods as baseline algorithms for comparison with FedEFC, evaluating its effectiveness in mitigating noisy labels and handling heterogeneous FL settings.
FL algorithms against noisy dataset: Several research efforts have explored ways to resolve the noise labels present in local datasets of FL. In [24], FedCorr is designed to cope with noisy data challenges. On the other hand, the framework requires the assumption that certain clients have entirely clean data to ensure improvement. FedNoRo [23] handles class imbalance and heterogeneous label noise in FL, but it requires accurate identification of noisy clients, which may be challenging in practice. Although RoFL [25] effectively leverages similarity-based learning to mitigate the issue of asymmetric noisy labels, it requires additional information on feature data that may pose a potential risk to privacy and communication bottleneck. All considered approaches, including FedCorr, are compared with the proposed method to provide a comprehensive and fair evaluation.
Figure 1: Overview of FedEFC framework. The scheme consists of two phases: (1) determining the prestopping point and (2) refining the loss correction. In Phase 1, the centralized server tracks client training accuracies to identify the prestopping point where model parameters are near-optimal. In Phase 2, each client updates its loss function using enhanced forward correction, guiding global model parameters toward their optimal configuration in the clean data space.
@@ 2.1 Data Allocation
Specifically, the Bernoulli probability $p$ determines whether examples of a specific label $i$ are allocated to client $k$. The Bernoulli distribution is represented by the indicator $\mathbb{I}(k,i)\sim p$, which takes the value $1$ if label $i$ is allocated to client $k$, and 0 otherwise. Once the Bernoulli distribution is established, the number of examples assigned to clients with $\mathbb{I}(\cdot,i)=1$ follows a Dirichlet distribution parameterized by $\alpha_{\text{dir}}>0$. Therefore, the degree of non-IID data allocation is controlled by the parameters $p$ and $\alpha_{\text{dir}}$.
@@ 2.2 Noise Generation
A symmetric noise flip more closely resembles real-world scenarios than the symmetric case when generating synthetic noise. To generate asymmetric noise, the sparsity mentioned in [16], [21] is utilized. The amount of noise $\rho$ and sparsity $\zeta$ need to be considered for noise generation. $\rho$ is the fraction of samples whose labels are flipped while $\zeta$ represents the proportion of uncorrupted labels, excluding the true label. For example, if all labels are flipped, $\zeta=0$. In contrast, $\zeta=1$ indicates that the dataset contains only clean labels. Thus, high sparsity represents greater imbalance in noisy labels.
@@ 3 Proposed Method
In this section, we propose FedEFC for mitigating the impact of noisy labels in non-IID FL. FedEFC consists of two phases: (1) determining a prestopping point and (2) applying loss correction. In Phase 1, the centralized server estimates the accuracy of the global model by aggregating the training accuracies of local clients. During Phase 1, each client transmits its measured accuracy to the centralized server, which monitors accuracy variations to determine the prestopping point. Once this point is reached, the server notifies the clients that Phase 1 is complete and that no further accuracy updates will be transmitted in subsequent rounds. In Phase 2, participating clients generate a noise transition matrix by analyzing the allocated dataset and apply loss correction. Further details are provided below.
@@ 3.1 Finding the Prestopping Point
where $A_{k}(w_{g}^{t})$ is training accuracy of client $k$ with the global model based on its own dataset. The training set contains both clean and noisy labels, which differs from the original heuristic validation approach that relies on a clean set [20]. A threshold $\gamma_{\text{thr}}$ is defined to quantify the instability in learning. In addition, a patience parameter $\tau_{p}$ is employed to track previous accuracy history. The accuracy measured in the current round, $A(t)$, is compared to the highest observed accuracy $A_{\text{max}}$, where $A_{\text{max}}=\max_{i<t}A(i)$ with $A_{\text{max}}$ initialized to 0. If $A(t)<A_{\text{max}}$, the patience parameter $\tau_{p}$ is incremented by 1; otherwise, $\tau_{p}$ is reset to 0 and $A_{\text{max}}$ is updated to $A(t)$. When the accuracy does not improve for $\gamma_{\text{thr}}$ consecutive rounds (i.e., $\tau_{p}=\gamma_{\text{thr}}$), the current round is regarded as the prestopping point, denoted as $T_{e}$. At the end of Phase 1, the centralized server instructs the clients to apply loss correction for every round after $T_{e}$. Empirical verification of the estimated accuracy, aggregated from each client’s training performance, is provided in the left panel of Fig. 2. In this panel, the estimated accuracy exhibits a sudden decline, attributed to the presence of noisy labels.
Figure 2: Left: Test accuracy and estimated accuracy used to determine the prestopping point. Right: Cosine similarity between the real noise matrix and the noise matrices estimated by the pretrained model and the global model in training.
@@ 3.2.1 Generating the Noise Transition Matrix
Each client acquires the noise transition matrix as follows. For client $k$, the global model after round $t=T_{e}$ predicts the label for input $\bm{x}^{k}$ as
$j=\arg\max_{j^{\prime}\in\mathcal{C}}\hat{p}_{k}(\tilde{y}^{k}=j^{\prime}|\bm{x}^{k},w^{t}_{g}),$ (3)
where $\hat{p}_{k}(\cdot)$ denotes an estimate of the conditional probability for client $k$. We define the set of examples with true label $y^{k}=j$ and observed label $\tilde{y}^{k}=i$ as
@@
$\quad\quad=\Bigl\{\bm{x}^{k}\in\mathbf{X}_{\tilde{y}^{k}=i}:\hat{p}_{k}(\tilde{y}^{k}=j|\bm{x}^{k};w^{t}_{g})\geq\tau_{j}\Bigr\},$ (4)
where the set $\mathbf{X}_{\tilde{y}^{k}=i}$ represents all data samples with observed label $i$ in the dataset $\mathcal{D}_{k}$ of client $k$. The example set $\mathbf{X}_{\tilde{y}^{k}=i}$ satisfies $\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\subseteq\mathbf{X}_{\tilde{y}^{k}=i}\subseteq\mathcal{D}_{k}$. The threshold $\tau_{j}$ serves as the criterion for counting examples and is defined as
$\tau_{j}=\frac{1}{\left|\mathbf{X}_{\tilde{y}^{k}=j}\right|}\sum_{\bm{x}^{k}\in\mathbf{X}_{\tilde{y}^{k}=j}}\hat{p}_{k}(\tilde{y}^{k}=j|\bm{x}^{k};w^{t}_{g}),$ (5)
Threshold $\tau_{j}$ is the mean estimation probabilities for data labeled as $j$. Using the example set $\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}$, the count matrix for client $k$ is defined as $\bm{C}_{\tilde{y}^{k}=i,y^{k}=j}=\left|\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\right|$. For simplicity, the notation is rewritten as $\bm{C}_{\tilde{y}^{k}=i,y^{k}=j}\coloneqq\bm{C}^{k}_{\tilde{y}=i,y=j},\mathbf{X}_{\tilde{y}^{k}=i,y^{k}=j}\coloneqq\mathbf{X}^{k}_{\tilde{y}=i,y=j}.$ The count matrix is constructed to enumerate the number of examples associated with each label pair, with columns representing true labels and rows indicating observed labels. Fig. 3 illustrates an example of count matrix construction. The noise transition matrix of client $k$ is computed as
@@
$\bm{Q}^{k}_{\tilde{y}=i|y=j}$ represents the $(i,j)$th entry of the noise transition matrix $\bm{Q}^{k}_{{\tilde{y}}|{y}}$. This element corresponds to an estimate of the conditional probability $p(\tilde{y}^{k}=i|y^{k}=j)$, which indicates the probability that the observed noisy label $\tilde{y}$ is $i$ given that the true label $y$ is $j$. Here, we assume that label noise is sample-independent, so $p(\tilde{y}^{k}=i|y^{k}=j,\bm{x}^{k})=p(\tilde{y}^{k}=i|y^{k}=j)$. Consequently, each client generates its own noise transition matrix $\bm{Q}^{k}_{{\tilde{y}}|{y}}$.
@@ 3.2.2 Applying Loss Correction
Participating clients apply loss correction based on the derived matrix $\bm{Q}_{{\tilde{y}}|{y}}$, following a similar approach to [17], where loss functions are modified. Regarding client $k$ in round $t$, the per-sample loss function $\mathcal{L}_{k}(\cdot)$ is given by $\mathcal{L}_{k}(\bm{x},y;{w}_{k}^{t})$. When a client trains with a DNN and applies the softmax function $\sigma_{k}(\cdot)$ in the final layer, the estimated conditional probability that the observed label is $\tilde{y}^{k}=i$ given input $\bm{x}^{k}$, using the client’s local model weights $w=w_{k}^{t}$ as our estimator parameterization, is
@@
$\bm{h}_{i}^{k}$ denotes the output of the DNN at the final layer corresponding to label $i$. The per-sample loss function $\mathcal{L}_{k}$ can be rewritten using cross-entropy as
$\mathcal{L}_{k}(\bm{x}^{k},i;w_{k}^{t})=-\log\hat{p}_{k}(\tilde{y}=i|\bm{x}^{k};w_{k}^{t}).$ (9)
To streamline notation, we replace $\bm{x}^{k}$, $y^{k}$ and $\tilde{y}^{k}$ with $\bm{x}$, $y$ and $\tilde{y}$, respectively. Let $\hat{p}_{k}(y=j|\bm{x};w)$ be an estimator for the conditional probability that the true label is $y=j$, given input $\bm{x}$, where this estimator is parametrized by $w$. Since we assumed $p(\tilde{y}|y,\bm{x})=p(\tilde{y}|y)$, it follows from Bayes’ rule that $p(\tilde{y}=i,y=j|\bm{x})=p(\tilde{y}=i|y=j)\,p(y=j|\bm{x})$ which implies
$p(\tilde{y}=i|\bm{x})=\sum_{j\in\mathcal{C}}p(\tilde{y}=i|y=j)\,p(y=j|\bm{x}).$ (10)
Thus, using clients local model weights $w=w_{k}^{t}$ as our estimator parametrization, and using the noise transition matrix entry $Q^{k}_{\tilde{y}=i|y=j}$ as our estimate for $p(\tilde{y}=i|y=j)$, we shall define the estimator
$\!\!\!\!\hat{p}_{k}({y}=i|\bm{x};w_{k}^{t}):=\sum_{j\in\mathcal{C}}Q^{k}_{\tilde{y}=i|y=j}\cdot\hat{p}_{k}({y}=j|\bm{x};w_{k}^{t}),$ (11)
and use this estimator to introduce an updated per-sample loss function $\mathcal{L}^{\textbf{F}}_{k}(\cdot)$, defined by
$\mathcal{L}^{\textbf{F}}_{k}(\bm{x}^{k},i;w_{k}^{t}):=-\log\hat{p}_{k}(y=i|\bm{x};w_{k}^{t}).$ (12)
Clients then begin transmitting their learning parameters, updated via loss correction, to the centralized server. The entire process is summarized in Algorithm 1.
@@
4:Initialize parameters $\tau_{p}=0$ and $A_{\text{max}}=0$
5:for $t=0$ to $T$ do
6:  $\mathcal{S}_{t}\leftarrow\text{Sample a subset of clients from }\{1,2,\dots,N\}$
7:  for each client $k\in\mathcal{S}_{t}$ in parallel do
8:   $\left(w_{k}^{t+1},A_{k}(w_{g}^{t})\right)\leftarrow\text{Client update}(k,w_{g}^{t})$
9:   Compute $A(t)$ via Eq. (2)
10:   if $A(t)>A_{\text{max}}$ then
11:     $\tau_{p}\leftarrow 0$, $A_{\text{max}}\leftarrow A(t)$
12:   else
13:     $\tau_{p}\leftarrow\tau_{p}+1$
14:     if $\tau_{p}=\gamma_{\text{thr}}$ then
15:      $T_{e}\leftarrow t$
16:      break
17:     end if
18:   end if
@@
20:  $w_{g}^{t+1}\leftarrow\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}}{\sum_{u\in\mathcal{S}_{t}}n_{u}}w_{k}^{t+1}$
21:end for
22:Phase 2: Apply loss correction
23:for $t=T_{e}$ to $T$ do
24:  $\mathcal{S}_{t}\leftarrow\text{Sample a subset of clients from }\{1,2,\dots,N\}$
25:  for each client $k\in\mathcal{S}_{t}$ in parallel do
26:   Compute ${Q}^{k}_{y=j|\tilde{y}=i}$ via Eq. (6)
27:   Update $\mathcal{L}_{k}^{\textbf{F}}(\bm{x},y;w_{k}^{t})$ via Eq. (12)
28:   $w_{k}^{t+1}\leftarrow\text{Client update}(k,w_{g}^{t},\mathcal{L}_{k}^{\textbf{F}}(\cdot))$
29:  end for
30:  $w_{g}^{t}\leftarrow\sum_{k\in\mathcal{S}_{t}}\frac{n_{k}}{\sum_{u\in\mathcal{S}_{t}}n_{u}}w_{k}^{t}$
31:end for